OpenStudy (zenmo):

Solve the differential equation. y ' = 9x - y

OpenStudy (phi):

I would write it in standard form y' - y = 9x first find the homogenous solution : y' - y = 0 do you know how to do that ?

OpenStudy (zenmo):

No

OpenStudy (phi):

you assume the solution has the form y= e^(ax) sub that into the equation. what do you get ?

OpenStudy (zenmo):

\[y ' + e^{ax} = 0\] ?

OpenStudy (phi):

yes, but if y= e^ax you can find y'

OpenStudy (phi):

what is d/dx of e^(ax) ?

OpenStudy (zenmo):

\[ae^{ax}\]

OpenStudy (phi):

yes, put that in for y'

OpenStudy (zenmo):

\[ae^{ax}+e^{ax} = 0\] ?

OpenStudy (phi):

yes e^ax is not zero (except at plus or minus infinity depending on the sign of a) so we can divide both sides by e^ax then solve for a

OpenStudy (zenmo):

\[a + 1 = 0\] ?

OpenStudy (zenmo):

\[\frac{ ae^{ax} }{ e^{ax} }+\frac{ e^{ax} }{ e^{ax} } = 0\]

OpenStudy (phi):

yes a+1=0 is called the characteristic equation. people usually "cut to the chase" and write it down immediately: y' + y =0 becomes a+1 = 0 (another example: y'' + 2y' -3y would become a^2 +2a -3 = 0 ) anyway, the homogenous solution is y = c e^(ax) with a=1: y = c e^x

OpenStudy (phi):

next we need to find the "particular solution" i.e. the solution that solves y' + y = 9x

OpenStudy (phi):

one strategy is to *assume* the solution has the same form as the right-hand side, and form a linear combination of it and its derivatives what that means is we assume the solution is Ax (A is an unknown coefficient) and the derivative of x: 1, but scaled by another unknown coefficient in other words, we try y = Ax + B put that into the equation y' + y = 9x what do you get ?

OpenStudy (zenmo):

\[y' + Ax + B = 9x\]?

OpenStudy (phi):

yes for y. do the same for y' (first find d/dx of y)

OpenStudy (zenmo):

\[\frac{ dy }{ dx }(Ax+B) = A, y ' = A\]\[A + Ax+B = 9x\] ?

OpenStudy (phi):

yes, now match up terms Ax + (A+B) = 9 x + 0

OpenStudy (phi):

A matches with 9 A+B with 0

OpenStudy (zenmo):

Okay

OpenStudy (phi):

we know A, what is B?

OpenStudy (zenmo):

B = - 9? Since Ax/x = 9.

OpenStudy (phi):

Ax + (A+B) = 9x + 0 the coeff in front the x's must match up. so A=9 the constant A+B = 0 with A=9 we have 9+B= 0 B= -9

OpenStudy (phi):

that makes the particular solution y= 9x - 9 the homogenous solution was from a+1 = 0 a= -1 so y= c e^-x (I make it e^x up above, but that was wrong) the total solution is the sum of the homogenous and particular solutions

OpenStudy (phi):

\[ y = c e^{-x} + 9x - 9 \] where c is an unknown constant. If we have initial conditions we can find its value.

OpenStudy (zenmo):

\[y=ce^{-x}+9x-9\] would be the answer if no initial conditions are given?

OpenStudy (phi):

yes. we can check it in the original equation first find y'

OpenStudy (zenmo):

\[y' = -ce^{-x}+9\]

OpenStudy (phi):

the other side is 9x - y = 9x - c e^-x -9x + 9 or -c^-x +9 (the 9x - 9x cancel) which matches y' (the left side)

OpenStudy (zenmo):

Thanks for the walk-through! :)